accents: é now inline: blaé test now inline: bla étest now inline: bla é test real spaces xxx tildes accent õ horizontal 2 - is – horizontal 3 - is — ligatures: ffl fi fl Yp

now in verbatim (tt font) ligatures: ffl fi fl Yp
what about vertikal bars |
horizontal bar -
and tildes ~
< funny carrots >